OpenStudy (anonymous):

1. John picked a bad of oranges to give to his friends. To the first of his friends, he gave half of the oranges he had and another one beside. to his second friend he gave half of the remaining oranges and another one beside. By this time, John had one orange left. How many did he start with?

OpenStudy (blank ):

10 Oranges

OpenStudy (blank ):

10/2 = 5 5 - 1 = 4 4/2 = 2 2 - 1 = 1 10 in all

OpenStudy (anonymous):

Wow! Thank you so much!

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

Wow that's fancy. Lol.